# Reconciling Kaplan and Chinchilla Scaling Laws

### v2

@@ Abstract
Kaplan et al. [2020] (‘Kaplan’) and Hoffmann et al. [2022] (‘Chinchilla’) studied the scaling behavior of transformers trained on next-token language prediction. These studies produced different estimates for how the number of parameters ($N$) and training tokens ($D$) should be set to achieve the lowest possible loss for a given compute budget ($C$). Kaplan: $N_{\text{optimal}}\propto C^{0.73}$, Chinchilla: $N_{\text{optimal}}\propto C^{0.50}$. This paper finds that much of this discrepancy can be attributed to Kaplan counting non-embedding rather than total parameters, combined with their analysis being performed at small scale. Simulating the Chinchilla study under these conditions produces biased scaling coefficients close to Kaplan’s. Hence, this paper reaffirms Chinchilla’s scaling coefficients, by explaining the primary cause of Kaplan’s original overestimation. As a second contribution, the paper explains differences in the reported relationships between loss and compute. These findings lead us to recommend that future scaling studies use total parameters and compute.  [^1] Code for analysis: https://github.com/TeaPearce/Reconciling_Kaplan_Chinchilla_Scaling_Laws
@@ 1 Introduction
Kaplan et al. [2020] (‘Kaplan’) and Hoffmann et al. [2022] (‘Chinchilla’) provided two influential studies measuring the impact of scale in large language models (LLMs). Both informed large-scale efforts on how to trade off model parameters ($N$) and data size ($D$) for a given compute budget ($C$), but with conflicting advice. Kaplan’s finding that $N_{\text{optimal}}\propto C^{0.73},D_{\text{optimal}}\propto C^{0.27}$ led to the conclusion that “big models may be more important than big data”, and LLMs trained in the ensuing years committed more resources to model size and less to data size. The subsequent Chinchilla study found $N_{\text{optimal}}\propto C^{0.50},D_{\text{optimal}}\propto C^{0.50}$, leading to their main thesis “for many current LLMs, smaller models should have been trained on more tokens to achieve the most performant model”, sparking a trend towards LLMs of more modest model sizes being trained on more data.
What was the cause of the difference in these scaling coefficient estimates that led to vast amounts of compute (plus emissions and finances) being used inefficiently? There have been suggestions that differences could be explained by different optimization schemes [Hoffmann et al., 2022] or datasets [Bi et al., 2024]. This note finds these suggestions incomplete, and offers a simple alternative explanation; much of the discrepancy can be attributed to Kaplan counting non-embedding rather than total parameters, combined with their analysis being performed at small scale.
@@
  • We develop an analytical approach comparing the scaling relationships reported in Kaplan and Chinchilla.
  • This approach finds that Kaplan’s reported relationship is locally consistent with Chinchilla’s, if non-embedding parameters are used, and at smaller scale.
  • Secondly, we study the reported relationships between compute and loss. Again non-embedding parameters and smaller scale models are the cause of Kaplan’s biased estimate, combined with the absence of an offset term in their compute-loss form.
  • We recommend that going forward the scaling community measures total parameters and compute, and uses an offset in the compute-loss relationship.
2 Set Up
Kaplan studied relationships in terms of non-embedding parameters ($N_{\setminus E}$) and non-embedding compute ($C_{\setminus E}$), excluding the linear layers embedding the vocabulary and position indices ($N_{E}$). By contrast, Chinchilla studied total parameters ($N_{T}$) and total compute ($C_{T}$). Define,
@@ 2 Set Up
where $d$ is dimension of the transformer residual stream, $v$ is vocab size, $h$ is context length (only included when positional embeddings are learned). Using the common approximation for FLOPs $C=6ND$, we define total and non-embedding compute,
@@
The estimated scaling coefficients can be more precisely written (using to signify ‘optimal’),
@@
Our approach uses information and data from the Chinchilla and Kaplan studies to estimate the scaling laws that would emerge if the Chinchilla relationship had been expressed in terms of $N_{\setminus E}$ $C_{\setminus E}$, and this had been done at the smaller model sizes used in Kaplan, as summarized in Figure 1.
@@
Finally we empirically verify our analysis in Section 4, by training a set of language models at tiny scale and conducting scaling law analyses under various settings. We find that simply changing the basis $N_{T}$ to $N_{\setminus E}$ produces coefficients inline with Chinchilla and Kaplan respectively, while multiple token budgets and decay schedules does not.
@@ 3 Analysis: Compute-Parameter Scaling Coefficient
where $l$ is number of layers. Whilst Kaplan do not list their model configurations, they do study varying aspect ratio $A=d/l$ for a fixed size model. They find that models of a given size perform similarly over a range of aspect ratios, and range is not affected by model scale (their Figure 5). Hence, we could assume a sizing scheme with fixed aspect ratio ($A\approx 40$ appears sensible from their plots). Assuming this sizing allows us to state,
@@
Chinchilla perspective. We empirically fit a function $N_{T}=N_{\setminus E}+\omega N_{\setminus E}^{\delta}$ (note the learnable exponent) to the Chinchilla model configurations listed in Table A9 of Hoffmann et al. [2022] for a range of $N_{T}$ (44M to 16B). We compute $N_{E}$ from Eq. 2, using the reported vocab size of 32,000, but ignore the context length 2,048 since Chinchilla used non-learnable position embeddings (though their inclusion effects coefficients only slightly).
@@
In order to analytically study their scaling relationship, we need an analytical from of loss. A functional form central to the Chinchilla study is
$L(N_{T},D)$ $=\frac{N_{c}}{N_{T}^{\alpha}}+\frac{D_{c}}{D^{\beta}}+E,$ (12)
$L(N_{T},C_{T})$ $=\frac{N_{c}}{N_{T}^{\alpha}}+\frac{D_{c}}{(C_{T}/6N_{T})^{\beta}}+E,$ (13)
where $N_{c},D_{c},\alpha,\beta>0$ are constants, $D$ is number of training tokens, $E$ is the “unpredictable” loss of the sequence. By differentiating and setting to zero, then rearranging in terms of $N_{T}$ we find
@@
To find the relationship between $N^{*}_{\setminus E}$ $C_{\setminus E}$ we take the derivative of Eq. 15, set to zero and rearrange,
@@
Figure 3 plots Eq. 16 and 17, using coefficients for $\alpha,\beta,N_{c},D_{c}$ from the Epoch AI specification (see later), and $\omega=47491$. There are three phases.
@@
Figure 3: Visualization of Eq. 16 & 17, using the Epoch AI specification.
@@
Figure 4 shows the synthetic training curves generated. We simulated 20 models with $N_{\setminus E}$ ranging from 790 parameters to 1.58B (Kaplan reports using model sizes “ranging in size from 768 to 1.5 billion non-embedding parameters”). For other constants in Eq. 15, there are two options. Whilst the original Chinchilla study reported finding constants of $N_{c}=406.4,Dc=410.7,\alpha=0.3392,\beta=0.2849,E=1.6934\implies N^{*}_{T}\propto C_{T}^{0.46}$ (‘Chinchilla specification’), Besiroglu et al. [2024] conducted a re-analysis, arguing that the following set of constants were more accurate $N_{c}=482.01,Dc=2085.43,\alpha=0.3478,\beta=0.3658,E=1.8172\implies N^{*}_{T}\propto C_{T}^{0.51}$ (‘Epoch AI specification’). Our figures adopt the Epoch AI specification and $\omega=47491$, though we report final results for the Chinchilla specification also.
@@
which are close to the Kaplan coefficient of 0.73.
Figure 4: Synthetic training curves from Eq. 15 using the Epoch AI specification, across 20 logarithmically-spaced models matching Kaplan’s size range. Left in terms of training tokens, right in terms of compute.
Figure 5: Using synthetic training curves generated in Figure 4, we empirically fit the frontier of compute efficient models using Chinchilla’s Method 1. This gives our main result; synthetic training curves generated from Chinchilla’s study, produce a local scaling coefficient ${N^{*}_{\setminus E}}\propto C_{\setminus E}^{0.78}$, close to Kaplan’s coefficient of 0.73. The analytical function from Eq. 16 is also verified.
@@ 4 Experiments: Compute-Parameter Scaling Coefficient
Result 2. Table 1 shows that optimization scheme has a smaller impact on scaling coefficients than switching from $N_{T}$ to $N_{\setminus E}$. Using a single set of models with no annealing (scheme 2) produces the same coefficients as using the more computationally expensive scheme 4. Counter to Chinchilla’s comment that moving from Kaplan’s scheme 3 to scheme 4 would reduce the scaling coefficient, our experiment suggests the opposite is the case, increasing from 0.46 to 0.49. This might explain our slight overestimation of the scaling coefficients in Eq. 52 & 61.
@@ 5 Analysis: Compute-Loss Scaling Coefficient
As well as using their analysis to calculate the compute-optimal parameter coefficient, Kaplan and Chinchilla also described the relationship between compute and loss, assuming parameters were scaled optimally. This optimal loss was again given in terms of non-embedding compute by Kaplan, and total compute by Chinchilla,
@@
Kaplan compute-loss fit: $L^{*}_{\setminus E}\propto C_{\setminus E}^{0.057}$ (39)
@@
Chinchilla compute-loss fit, Epoch AI spec: $L^{*}_{T}-E\propto C_{T}^{0.178}$ (41)
Chinchilla spec: $L^{*}_{T}-E\propto C_{T}^{0.155}$ (42)
@@
Our analysis follows the four-step approach from Section 3. We can directly reuse Step 1 & 2, while Step 3 & 4 are now modified to study the relationship between optimal loss and compute, rather than optimal parameters and compute as previously.
@@
Figure 7 visualizes the fit of the Chinchilla compute-loss form vs. the Kaplan compute-loss form, when one counts total compute vs. non-embedding compute. We see that Kaplan’s form provides a good fit of the data in the non-embedding compute plot, over the range of model sizes they considered. We speculate that this might be the motivation for Kaplan’s selection of this simpler compute-loss form.
@@ 6 Related work
Following early works formalizing how language models improve with parameters, data, and training compute [Rosenfeld et al., 2019, Kaplan et al., 2020, Hoffmann et al., 2022], there has been investigation into whether these scaling laws arise in other domains [Henighan et al., 2020], and to explain their existence from a theoretical standpoint [Hutter, 2021, Maloney et al., 2022, Bahri et al., 2024].
Closer in spirit to our paper, are several concurrent works that have investigated various design decisions in scaling law analyses. Su et al. [2024] revisit the methodology used to find scaling coefficients. Hägele et al. [2024] found that multiple independent cosine schedules could be reproduced more efficiently through a constant learning rate with multiple short decays, or stochastic weight averaging. Our finding is subtly different; a simple fixed learning rate will recover very similar compute-parameter scaling coefficients as multiple cosine schedules. Bi et al. [2024] study the effect of various hyperparameters on scaling. They observe that different text datasets produce slightly different optimal coefficients, with ‘cleaner’ data leading to more parameter-hungry scaling behavior, which they speculate could partially explain the difference between Kaplan and Chinchilla coefficients.
Porian et al. [2024] provide a concurrent work with the same objective as our paper – explaining the differences between the Kaplan and Chinchilla coefficients. Through a set of large-scale experiments reproducing Kaplan’s study, they determine that responsibility for the discrepancy can be attributed, in decreasing order of significance, to; 1) Kaplan counting non-embedding rather than total compute. 2) Kaplan using a fixed-length warmup period that was too long for smaller models, making them appear less efficient. 3) Kaplan not fully tuning optimization hyperparameters. We see these findings as complimentary to our own. We have been able to identify the primary ‘first-order’ reason using only information that was publicly available in the two papers, with a fully analytical approach. (Tiny-scale experiments were run post-hoc as verification.) This illustrates the promise of applying mathematical approaches to the empirical science of scaling.

### v3

@@ Abstract
Kaplan et al. (2020) (‘Kaplan’) and Hoffmann et al. (2022) (‘Chinchilla’) studied the scaling behavior of transformers trained on next-token language prediction. These studies produced different estimates for how the number of parameters ($N$) and training tokens ($D$) should be set to achieve the lowest possible loss for a given compute budget ($C$). Kaplan: $N_{\text{optimal}}\propto C^{0.73}$, Chinchilla: $N_{\text{optimal}}\propto C^{0.50}$. This paper finds that much of this discrepancy can be attributed to Kaplan counting non-embedding rather than total parameters, combined with their analysis being performed at small scale. Simulating the Chinchilla study under these conditions produces biased scaling coefficients close to Kaplan’s. Hence, this paper reaffirms Chinchilla’s scaling coefficients, by explaining the primary cause of Kaplan’s original overestimation. As a second contribution, the paper explains differences in the reported relationships between loss and compute. These findings lead us to recommend that future scaling studies use total parameters and compute.  [^1] Code for analysis: https://github.com/TeaPearce/Reconciling_Kaplan_Chinchilla_Scaling_Laws
@@ 1 Introduction
Kaplan et al. (2020) (‘Kaplan’) and Hoffmann et al. (2022) (‘Chinchilla’) provided two influential studies measuring the impact of scale in large language models (LLMs). Both informed large-scale efforts on how to trade off model parameters ($N$) and training tokens ($D$) for a given compute budget ($C$), but with conflicting advice. Kaplan’s finding that $N_{\text{optimal}}\propto C^{0.73},D_{\text{optimal}}\propto C^{0.27}$ led to the conclusion that “big models may be more important than big data”, and LLMs trained in the ensuing years committed more resources to model size and less to data size. The subsequent Chinchilla study found $N_{\text{optimal}}\propto C^{0.50},D_{\text{optimal}}\propto C^{0.50}$, leading to their main thesis “for many current LLMs, smaller models should have been trained on more tokens to achieve the most performant model”, sparking a trend towards LLMs of more modest model sizes being trained on more data.
What was the cause of the difference in these scaling coefficient estimates that led to vast amounts of compute (plus emissions and finances) being used inefficiently? There have been suggestions that differences could be explained by different optimization schemes (Hoffmann et al., 2022) or datasets (Bi et al., 2024). This note finds these suggestions incomplete, and offers a simple alternative explanation; much of the discrepancy can be attributed to Kaplan counting non-embedding rather than total parameters, combined with their analysis being performed at small scale.
@@
  • We develop an analytical approach comparing the scaling relationships reported in Kaplan and Chinchilla (Section 3). This approach finds that Kaplan’s reported relationship is locally consistent with Chinchilla’s, if non-embedding parameters are used, and at smaller scale.
  • Secondly, we study the reported relationships between compute and loss (Section 5). Again non-embedding parameters and smaller scale models are the cause of Kaplan’s biased estimate, combined with the absence of an offset term in their compute-loss form.
  • We recommend that going forward the scaling community measures total parameters and total compute, and uses an offset in the compute-loss relationship.
2 Preliminaries
This section introduces some background knowledge and definitions (Section 2.1), outlines the analysis approach for our main result (Section 2.2), and documents assumptions (Section 2.3).
2.1 Set Up
Neural scaling laws.
Kaplan et al. (2020) & Hoffmann et al. (2022) investigated empirically modeling the relationships between the number of parameters ($N$), training tokens ($D$), training compute ($C$) and loss ($L$) in transformers for language modeling. The main functional form of relationship considered was a power law, $y=ax^{b}$, which is widely used throughout the sciences to describe a relationship between two quantities ($x$ $y$) that holds over many orders of magnitude, e.g. (Kello et al., 2010).
Definition of $N$ $C$. One difference between the two studies is the definition of $N$ $C$. Kaplan studied relationships in terms of non-embedding parameters ($N_{\setminus E}$) and non-embedding compute ($C_{\setminus E}$), excluding the contributions of the embedding layers for the vocabulary and position indices ($N_{E}$). By contrast, Chinchilla studied total parameters ($N_{T}$) and total compute ($C_{T}$). We define,
@@ 2.1 Set Up
where $d$ is the dimension of the transformer residual stream, $v$ is vocab size, $h$ is context length (only included when positional embeddings are learned). Using the common approximation for training compute FLOPs $C=6ND$ (a factor of 6 covers a forward and backward pass), we define total and non-embedding compute,
@@
Compute optimality. This definition of compute $C=6ND$ suggests a direct trade off between parameters and training tokens for a given compute budget. The two studies focus on ‘compute optimal’ configurations of these quantities. That is, for a given compute budget, the parameters and training tokens that lead to the lowest possible loss. For total parameters this is written (using to signify ‘optimal’),
Given this notation, the estimated scaling coefficients can be more precisely written,
@@
Functional form. An important functional form relating $N_{T}$, $D$, $L$ used in Chinchilla is given by,
$L(N_{T},D)$ $=\frac{N_{c}}{N_{T}^{\alpha}}+\frac{D_{c}}{D^{\beta}}+E,$ (8)
where $N_{c},D_{c},\alpha,\beta>0$ are empirically fitted constants and $E$ is the ‘irreducible’ loss inherent in language. This form conveniently gives rise to power law relationships; $N^{*}_{T}\propto C_{T}^{a}$ with $a=\frac{\beta}{\alpha+\beta}$, $D^{*}_{T}\propto C_{T}^{b}$ with $b=\frac{\alpha}{\alpha+\beta}$, and $L^{*}_{T}-E\propto C_{T}^{-\gamma}$ with $\gamma=\frac{\alpha\beta}{\alpha+\beta}$.
There are two possible specifications from the constants in Eq. 8 – those originally reported in Chinchilla, and those reported in a re-analysis conducted by Besiroglu et al. (2024) claiming to correct minor errors in the fitting procedure. Our work reports results with both specifications.
Chinchilla specification: $N_{c}=406.4,Dc=410.7,\alpha=0.3392,\beta=0.2849,E=1.693\implies N^{*}_{T}\propto C_{T}^{0.46},$ (9)
Epoch AI specification: $N_{c}=482.0,Dc=2085.43,\alpha=0.3478,\beta=0.3658,E=1.817\implies N^{*}_{T}\propto C_{T}^{0.51}.$ (10)
2.2 Analysis Overview
Our analysis uses information and data from the Chinchilla and Kaplan studies to estimate the scaling laws that would emerge if the Chinchilla relationship had been expressed in terms of $N_{\setminus E}$ $C_{\setminus E}$, and this had been done over the smaller model sizes used in Kaplan, as summarized in Figure 1.
@@ 2.2 Analysis Overview
Section 4 experimentally verifies our analysis by training a set of language models at tiny scale and conducting scaling law analyses under various settings. Simply changing the basis $N_{T}$ to $N_{\setminus E}$ produces coefficients inline with Chinchilla and Kaplan respectively, while multiple token budgets and decay schedules does not.
@@
2.3 Assumptions
For transparency, we list the assumptions and approximations made in our analysis.
  • We assume $C_{\setminus E}=6N_{\setminus E}D$ and $C_{T}=6N_{T}D$.
  • We assume a fixed functional form between total and non-embedding parameters in Eq. 11, and fit $\omega$ empirically using Chinchilla model configurations.
  • We assume a fixed functional form between loss, total parameters and training data given by Eq. 8. We report results using both the Chinchilla (Eq. 9) and Epoch AI (Eq. 10) fitted constants.
  • We approximate Kaplan’s models with 20 logarithmically spaced model sizes from 0.79k to 1.58B non-embedding parameters.
@@ 3 Analysis: Compute-Parameter Scaling Coefficient
where $l$ is number of layers. Whilst Kaplan do not list their model configurations, they do study varying aspect ratio $A=d/l$ for a fixed size model. They find that models of a given size perform similarly over a range of aspect ratios, and this is not affected by model scale (their Figure 5). Hence, we could assume a sizing scheme with fixed aspect ratio ($A\approx 40$ appears sensible from their plots). Assuming this sizing allows us to state (with $l=d/A$ in Eq. 12),
@@
Chinchilla perspective. We empirically fit a function $N_{T}=N_{\setminus E}+\omega N_{\setminus E}^{\delta}$ (note the learnable exponent) to the Chinchilla model configurations listed in Table A9 of Hoffmann et al. (2022) for a range of $N_{T}$ (44M to 16B). We compute $N_{E}$ from Eq. 2, using the reported vocab size of 32,000, but ignore the context length 2,048 since Chinchilla used non-learnable position embeddings (though their inclusion effects coefficients only slightly).
@@
In order to analytically study their scaling relationship, we need an analytical form of loss, for which we use the functional form in the Chinchilla study. Taking Eq. 8 and using $C_{T}=6N_{T}D$,
$L(N_{T},C_{T})$ $=\frac{N_{c}}{N_{T}^{\alpha}}+\frac{D_{c}}{(C_{T}/6N_{T})^{\beta}}+E.$ (16)
By differentiating Eq. 16 wrt $N_{T}$ and setting to zero, then rearranging in terms of $N_{T}$ we find
@@
To find the relationship between $N^{*}_{\setminus E}$ $C_{\setminus E}$ we take the derivative of Eq. 18 wrt $N_{\setminus E}$, set to zero and rearrange,
@@
Working is given in Appendix A.1. Figure 3 plots Eq. 19 & 20, using coefficients for $\alpha,\beta,N_{c},D_{c}$ from the Epoch AI specification, and $\omega=47491$. There are three phases.
@@
Figure 3: Visualization of Eq. 19 & 20, using the Epoch AI specification.
@@
Figure 4 shows the synthetic training curves generated. We simulated 20 models with $N_{\setminus E}$ ranging from 790 parameters to 1.58B (Kaplan reports using model sizes “ranging in size from 768 to 1.5 billion non-embedding parameters”). For other constants in Eq. 18, we adopt the Epoch AI specification (Eq. 10) and $\omega=47491$, though we report final results for the Chinchilla specification (Eq. 9) also.
@@
which are close to the Kaplan coefficient of 0.73. Hence, this shows that the Chinchilla coefficient is roughly consistent with Kaplan’s coefficient, given these two adjustments. This constitutes the paper’s main result, reconciling these two apparently conflicting results.
Figure 4: Synthetic training curves from Eq. 18 fitted to the Chinchilla data using the Epoch AI specification. Curves are generated for 20 logarithmically-spaced models matching Kaplan’s size range. Left in terms of training tokens, right in terms of non-embedding compute, as used in the Kaplan study. Hence the right plot can be viewed as Chinchilla’s loss curves, adjusted to match Kaplan’s model sizes and non-embedding measurements.
Figure 5: Using the synthetic training curves generated in Figure 4, we empirically fit the frontier of compute efficient models using Chinchilla’s Method 1. This gives our main result; synthetic training curves generated from Chinchilla’s study, adjusted for model sizes and non-embedding compute, produce a local scaling coefficient ${N^{*}_{\setminus E}}\propto C_{\setminus E}^{0.78}$, close to Kaplan’s reported coefficient of 0.73. The analytical function from Eq. 19 is also verified.
@@ 4 Experiments: Compute-Parameter Scaling Coefficient
Result 2. Table 1 shows that optimization scheme has a smaller impact on scaling coefficients than switching from $N_{T}$ to $N_{\setminus E}$. Using a single set of models with no annealing (scheme 2) produces the same coefficients as using the more computationally expensive scheme 4. Counter to Chinchilla’s comment that moving from Kaplan’s scheme 3 to scheme 4 would reduce the scaling coefficient, our experiment suggests the opposite is the case, increasing from 0.46 to 0.49. This might explain our slight overestimation of the scaling coefficients in Eq. 29 & 38.
@@ 5 Analysis: Compute-Loss Scaling Coefficient
As well as analyzing the compute-optimal parameter scaling, Kaplan and Chinchilla also described the scaling relationship between compute and loss, assuming parameters were scaled optimally. This optimal loss was again given in terms of non-embedding compute by Kaplan, and total compute by Chinchilla,
@@
Kaplan compute-loss fit: $L^{*}_{\setminus E}\propto C_{\setminus E}^{-0.057}$ (42)
@@
Chinchilla compute-loss fit, Epoch AI spec: $L^{*}_{T}-E\propto C_{T}^{-0.178}$ (44)
Chinchilla spec: $L^{*}_{T}-E\propto C_{T}^{-0.155}$ (45)
@@
Our analysis follows the same four-step approach as in Section 3. We can directly reuse Steps 1 & 2, while Steps 3 & 4 are now modified to study the relationship between optimal loss and compute, rather than optimal parameters and compute as previously.
@@
Figure 7 visualizes the fit of the Chinchilla compute-loss form vs. the Kaplan compute-loss form, when one counts total compute vs. non-embedding compute. We see that Kaplan’s form provides a good fit of the data in the non-embedding compute plot at small scale, over the range of model sizes they considered. We speculate that this might be the motivation for Kaplan’s selection of this simpler compute-loss form.
@@ 6 Related work
Following early works formalizing how language models improve with parameters, data, and training compute (Rosenfeld et al., 2019; Kaplan et al., 2020; Hoffmann et al., 2022), there has been investigation into whether these scaling laws arise in other domains (Henighan et al., 2020), and to explain their existence from a theoretical standpoint (Hutter, 2021; Maloney et al., 2022; Bahri et al., 2024).
Closer in spirit to our paper are several concurrent works that have investigated the influence of various design decisions on scaling law analyses. Su et al. (2024) revisit the methodology used to find scaling coefficients. Hägele et al. (2024) found that multiple independent cosine schedules could be reproduced more efficiently through a constant learning rate with multiple short decays, or stochastic weight averaging. Our finding is subtly different; a simple fixed learning rate will recover very similar compute-parameter scaling coefficients as multiple cosine schedules. Bi et al. (2024) study the effect of various hyperparameters on scaling laws. They observe that different text datasets produce slightly different optimal coefficients, with ‘cleaner’ data leading to more parameter-hungry scaling behavior, which they speculate could partially explain the difference between Kaplan and Chinchilla coefficients.
Porian et al. (2024) provide a concurrent work with the same objective as our paper – explaining the differences between the Kaplan and Chinchilla coefficients. Through a set of large-scale experiments reproducing Kaplan’s study, they determine that responsibility for the discrepancy can be attributed, in decreasing order of significance, to; 1) Kaplan counting non-embedding rather than total compute. 2) Kaplan using a fixed-length warmup period that was too long for smaller models, making them appear less efficient. 3) Kaplan not fully tuning optimization hyperparameters. We see these findings as complimentary to our own. We have been able to identify the primary ‘first-order’ reason using only information that was publicly available in the two papers, with a fully analytical approach. (Tiny-scale experiments were run post-hoc as verification.) This illustrates the promise of applying mathematical approaches to the empirical science of scaling.